OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x + 3y = 4 x - 2y = -5 a. 2 b. -2 c. 1 d. -1

jhonyy9 (jhonyy9):

what method you need using ?

OpenStudy (anonymous):

any

OpenStudy (anonymous):

\[2x + 3y = 4 \\x - 2y = -5\] multiply the second equation by \(-2\) and add it to the first

OpenStudy (anonymous):

?

OpenStudy (anonymous):

-2(x-2y=-5)= -2x+4y=10

OpenStudy (anonymous):

Now add it to the other equation.

OpenStudy (anonymous):

i still dont understand /:

OpenStudy (anonymous):

2x+3y=4 -2x+4y=10 7y=14 y=2 Do you know what to do now?

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

No. That gives you the value of y. You asked for the value of x.

OpenStudy (anonymous):

To get the value of x, plug in the value of y into any of the first two equations and solve for x.

OpenStudy (anonymous):

i get it now! thanks

OpenStudy (anonymous):

No problem. Tell us what you get.

OpenStudy (anonymous):

x= -1

OpenStudy (anonymous):

Great!